# EnsemW2S: Can an Ensemble of LLMs be Leveraged to Obtain a Stronger LLM?

### v1

@@ Appendix D Important Notations
Appendix E Binary Classification Task
Algorithm 2 AdaBoost Freund & Schapire (1997)
Training Dataset $S=\{(x_{i},y_{i})\}_{i=1}^{m}\sim D^{m}$
$T=$ AdaBoost iterations
$\vec{D}_{1}(i)\leftarrow\frac{1}{m}\forall i\in[m]$
for $t\leftarrow 1$ to $T$  do
     $\alpha_{t}\leftarrow\frac{1}{2}\log{\frac{1-\epsilon_{t}}{\epsilon_{t}}}$
     $Z_{t}\leftarrow 2\sqrt{\epsilon_{t}(1-\epsilon_{t})}$
     $\vec{D}_{t+1}\leftarrow\frac{1}{Z_{t}}\vec{D}_{t}e^{-\alpha_{t}y_{i}h_{t}(x_{i})}$
     $g\leftarrow\sum_{t=1}^{T}\alpha_{t}h_{t}$
Return $h(x)=\operatorname{sign}(g)$
E.1 Detailed Results for Binary Classification Task with $\alpha$ and $Err_{t}^{Train}$ in Table 1 and Table 2
AdaBoost Weak Model Strong Model Weak-to-Strong Model $\alpha$ $Err_{t}^{Train}$
Model Name GPT-2 GPT-2 Medium
Baseline 0.610 0.665 0.590 0.455 0.287
With AdaBoost (T:02) 0.613 0.665 0.619 0.488 0.274
With AdaBoost (T:03) 0.614 0.665 0.609 0.463 0.284
With AdaBoost (T:04) 0.611 0.665 0.622 0.467 0.282
With AdaBoost (T:05) 0.623 0.665 0.640 0.448 0.290
With AdaBoost (T:06) 0.621 0.665 0.641 0.333 0.340
With AdaBoost (T:07) 0.646 0.665 0.638 0.433 0.300
With AdaBoost (T:08) 0.610 0.665 0.626 0.471 0.281
With AdaBoost (T:09) 0.634 0.665 0.619 0.463 0.284
With AdaBoost (T:10) 0.618 0.665 0.622 0.503 0.268
Model Name GPT-2 GPT-2 Large
Baseline 0.610 0.681 0.591 0.455 0.287
With AdaBoost (T:02) 0.613 0.681 0.657 0.488 0.274
With AdaBoost (T:03) 0.614 0.681 0.620 0.463 0.284
With AdaBoost (T:04) 0.611 0.681 0.629 0.467 0.282
With AdaBoost (T:05) 0.623 0.681 0.656 0.448 0.290
With AdaBoost (T:06) 0.621 0.681 0.650 0.333 0.340
With AdaBoost (T:07) 0.646 0.681 0.654 0.433 0.300
With AdaBoost (T:08) 0.610 0.681 0.633 0.471 0.281
With AdaBoost (T:09) 0.634 0.681 0.648 0.463 0.284
With AdaBoost (T:10) 0.618 0.681 0.652 0.503 0.268
Model Name GPT-2 GPT-2 XL
Baseline 0.607 0.707 0.620 0.455 0.287
With AdaBoost (T:02) 0.613 0.707 0.654 0.488 0.274
With AdaBoost (T:03) 0.614 0.707 0.628 0.463 0.284
With AdaBoost (T:04) 0.611 0.707 0.663 0.467 0.282
With AdaBoost (T:05) 0.623 0.707 0.645 0.448 0.290
With AdaBoost (T:06) 0.621 0.707 0.648 0.333 0.340
With AdaBoost (T:07) 0.646 0.707 0.649 0.433 0.300
With AdaBoost (T:08) 0.610 0.707 0.653 0.471 0.281
With AdaBoost (T:09) 0.634 0.707 0.657 0.463 0.284
With AdaBoost (T:10) 0.618 0.707 0.654 0.503 0.268
Model Name GPT-2 Qwen1.5-1.8B
Baseline 0.602 0.842 0.646 0.445 0.291
With AdaBoost (T:02) 0.599 0.842 0.683 0.500 0.269
With AdaBoost (T:03) 0.626 0.842 0.702 0.444 0.292
With AdaBoost (T:04) 0.611 0.842 0.723 0.400 0.310
With AdaBoost (T:05) 0.613 0.842 0.704 0.461 0.285
With AdaBoost (T:06) 0.613 0.842 0.734 0.417 0.303
With AdaBoost (T:07) 0.603 0.842 0.712 0.422 0.301
With AdaBoost (T:08) 0.608 0.842 0.717 0.319 0.346
With AdaBoost (T:09) 0.614 0.842 0.712 0.405 0.308
With AdaBoost (T:10) 0.606 0.842 0.712 0.360 0.328
Model Name GPT-2 Medium GPT-2 Large
Baseline 0.653 0.681 0.626 0.705 0.196
With AdaBoost (T:02) 0.656 0.681 0.643 0.624 0.223
With AdaBoost (T:03) 0.646 0.681 0.639 0.674 0.206
With AdaBoost (T:04) 0.663 0.681 0.664 0.645 0.216
With AdaBoost (T:05) 0.645 0.681 0.654 0.690 0.201
With AdaBoost (T:06) 0.652 0.681 0.667 0.619 0.225
With AdaBoost (T:07) 0.650 0.681 0.665 0.722 0.191
With AdaBoost (T:08) 0.657 0.681 0.685 0.733 0.187
With AdaBoost (T:09) 0.651 0.681 0.684 0.601 0.231
With AdaBoost (T:10) 0.648 0.681 0.666 0.682 0.203
Model Name GPT-2 Medium GPT-2 XL
Baseline 0.653 0.707 0.655 0.705 0.196
With AdaBoost (T:02) 0.656 0.707 0.651 0.624 0.223
With AdaBoost (T:03) 0.646 0.707 0.648 0.674 0.206
With AdaBoost (T:04) 0.663 0.707 0.675 0.645 0.216
With AdaBoost (T:05) 0.645 0.707 0.663 0.690 0.201
With AdaBoost (T:06) 0.652 0.707 0.682 0.619 0.225
With AdaBoost (T:07) 0.650 0.707 0.657 0.722 0.191
With AdaBoost (T:08) 0.657 0.707 0.673 0.733 0.187
With AdaBoost (T:09) 0.651 0.707 0.665 0.601 0.231
With AdaBoost (T:10) 0.648 0.707 0.687 0.682 0.203
Model Name GPT-2 Medium Qwen1.5-1.8B
Baseline 0.649 0.842 0.722 0.658 0.211
With AdaBoost (T:02) 0.649 0.842 0.742 0.626 0.222
With AdaBoost (T:03) 0.669 0.842 0.732 0.673 0.206
With AdaBoost (T:04) 0.649 0.842 0.757 0.662 0.210
With AdaBoost (T:05) 0.661 0.842 0.745 0.688 0.202
With AdaBoost (T:06) 0.655 0.842 0.735 0.722 0.191
With AdaBoost (T:07) 0.664 0.842 0.732 0.717 0.192
With AdaBoost (T:08) 0.664 0.842 0.741 0.718 0.192
With AdaBoost (T:09) 0.657 0.842 0.748 0.791 0.171
With AdaBoost (T:10) 0.667 0.842 0.737 0.671 0.207
Model Name GPT-2 Large GPT-2 XL
Baseline 0.673 0.707 0.682 1.675 0.034
With AdaBoost (T:02) 0.658 0.707 0.675 0.974 0.125
With AdaBoost (T:03) 0.671 0.707 0.687 1.091 0.101
With AdaBoost (T:04) 0.671 0.707 0.684 1.080 0.103
With AdaBoost (T:05) 0.668 0.707 0.687 1.033 0.112
With AdaBoost (T:06) 0.675 0.707 0.683 1.133 0.094
With AdaBoost (T:07) 0.669 0.707 0.688 1.083 0.103
With AdaBoost (T:08) 0.676 0.707 0.683 1.047 0.110
With AdaBoost (T:09) 0.678 0.707 0.682 1.085 0.103
With AdaBoost (T:10) 0.669 0.707 0.681 1.132 0.094
Model Name GPT-2 Large Qwen1.5-1.8B
Baseline 0.664 0.842 0.749 1.454 0.052
With AdaBoost (T:02) 0.670 0.842 0.717 0.971 0.126
With AdaBoost (T:03) 0.670 0.842 0.728 0.037 0.481
With AdaBoost (T:04) 0.677 0.842 0.727 1.128 0.095
With AdaBoost (T:05) 0.675 0.842 0.740 1.107 0.098
With AdaBoost (T:06) 0.677 0.842 0.737 0.979 0.124
With AdaBoost (T:07) 0.676 0.842 0.766 1.136 0.093
With AdaBoost (T:08) 0.680 0.842 0.741 1.103 0.099
With AdaBoost (T:09) 0.691 0.842 0.762 1.075 0.104
With AdaBoost (T:10) 0.683 0.842 0.755 1.052 0.109
Model Name GPT-2 XL Qwen1.5-1.8B
Baseline 0.673 0.842 0.733 0.564 0.244
With AdaBoost (T:02) 0.701 0.842 0.740 0.428 0.298
With AdaBoost (T:03) 0.702 0.842 0.753 0.383 0.317
With AdaBoost (T:04) 0.694 0.842 0.756 0.316 0.347
With AdaBoost (T:05) 0.704 0.842 0.759 0.260 0.373
With AdaBoost (T:06) 0.693 0.842 0.757 0.288 0.360
With AdaBoost (T:07) 0.708 0.842 0.755 0.277 0.365
With AdaBoost (T:08) 0.706 0.842 0.761 0.223 0.391
With AdaBoost (T:09) 0.700 0.842 0.748 0.252 0.377
With AdaBoost (T:10) 0.703 0.842 0.747 0.258 0.374
Table 1: This table shows weak to strong generalization using random data-splits for sciq dataset. We also study the impact of using ensemble learning methods like AdaBoost, which combines weak learners, for weak to strong training. Each model is trained for 3 epochs and uses an optimized learning rate.
AdaBoost Weak Model Strong Model Weak-to-Strong $\alpha$ $Err_{t}^{Train}$
Model Name GPT-2 GPT-2 Medium
Baseline 0.362 0.638 0.388 2.178 0.013
With AdaBoost (T:02) 0.356 0.638 0.382 1.790 0.027
With AdaBoost (T:03) 0.343 0.638 0.386 1.953 0.020
With AdaBoost (T:04) 0.361 0.638 0.385 2.014 0.018
With AdaBoost (T:05) 0.361 0.638 0.382 1.534 0.044
With AdaBoost (T:06) 0.365 0.638 0.393 1.588 0.040
With AdaBoost (T:07) 0.365 0.638 0.402 1.474 0.050
With AdaBoost (T:08) 0.369 0.638 0.404 1.478 0.049
With AdaBoost (T:09) 0.362 0.638 0.394 1.865 0.023
With AdaBoost (T:10) 0.364 0.638 0.394 1.267 0.074
Model Name GPT-2 GPT-2 Large
Baseline 0.362 0.597 0.385 2.178 0.013
With AdaBoost (T:02) 0.356 0.597 0.367 1.790 0.027
With AdaBoost (T:03) 0.343 0.597 0.383 1.953 0.020
With AdaBoost (T:04) 0.361 0.597 0.379 2.014 0.018
With AdaBoost (T:05) 0.361 0.597 0.387 1.534 0.044
With AdaBoost (T:06) 0.365 0.597 0.382 1.588 0.040
With AdaBoost (T:07) 0.365 0.597 0.388 1.474 0.050
With AdaBoost (T:08) 0.369 0.597 0.389 1.478 0.049
With AdaBoost (T:09) 0.362 0.597 0.393 1.865 0.023
With AdaBoost (T:10) 0.364 0.597 0.395 1.267 0.074
Model Name GPT-2 GPT-2 XL
Baseline 0.355 0.561 0.421 2.178 0.013
With AdaBoost (T:02) 0.356 0.561 0.409 1.791 0.027
With AdaBoost (T:03) 0.343 0.561 0.409 1.953 0.020
With AdaBoost (T:04) 0.361 0.561 0.407 2.014 0.018
With AdaBoost (T:05) 0.361 0.561 0.418 1.534 0.044
With AdaBoost (T:06) 0.365 0.561 0.409 1.588 0.040
With AdaBoost (T:07) 0.365 0.561 0.407 1.474 0.050
With AdaBoost (T:08) 0.369 0.561 0.413 1.478 0.049
With AdaBoost (T:09) 0.362 0.561 0.410 1.865 0.023
With AdaBoost (T:10) 0.364 0.561 0.409 1.267 0.074
Model Name GPT-2 Qwen1.5-1.8B
Baseline 0.364 0.760 0.407 2.178 0.013
With AdaBoost (T:02) 0.356 0.760 0.397 1.791 0.027
With AdaBoost (T:03) 0.343 0.760 0.393 1.953 0.020
With AdaBoost (T:04) 0.361 0.760 0.381 2.014 0.018
With AdaBoost (T:05) 0.361 0.760 0.390 1.534 0.044
With AdaBoost (T:06) 0.365 0.760 0.394 1.588 0.040
With AdaBoost (T:07) 0.365 0.760 0.390 1.474 0.050
With AdaBoost (T:08) 0.369 0.760 0.387 1.478 0.049
With AdaBoost (T:09) 0.362 0.760 0.402 1.865 0.023
With AdaBoost (T:10) 0.364 0.760 0.404 1.267 0.074
Model Name GPT-2 Medium GPT-2 Large
Baseline 0.391 0.597 0.420 1.511 0.046
With AdaBoost (T:02) 0.448 0.597 0.438 1.571 0.041
With AdaBoost (T:03) 0.426 0.597 0.405 1.483 0.049
With AdaBoost (T:04) 0.454 0.597 0.437 1.601 0.039
With AdaBoost (T:05) 0.448 0.597 0.428 1.334 0.065
With AdaBoost (T:06) 0.465 0.597 0.444 1.249 0.076
With AdaBoost (T:07) 0.449 0.597 0.453 1.460 0.051
With AdaBoost (T:08) 0.461 0.597 0.444 1.646 0.036
With AdaBoost (T:09) 0.449 0.597 0.433 1.453 0.052
With AdaBoost (T:10) 0.447 0.597 0.424 1.154 0.090
Model Name GPT-2 Medium GPT-2 XL
Baseline 0.392 0.561 0.440 1.510 0.047
With AdaBoost (T:02) 0.459 0.561 0.442 1.589 0.040
With AdaBoost (T:03) 0.420 0.561 0.435 1.669 0.034
With AdaBoost (T:04) 0.458 0.561 0.441 1.460 0.051
With AdaBoost (T:05) 0.424 0.561 0.431 1.393 0.058
With AdaBoost (T:06) 0.444 0.561 0.448 1.286 0.071
With AdaBoost (T:07) 0.419 0.561 0.436 1.429 0.054
With AdaBoost (T:08) 0.454 0.561 0.443 1.596 0.039
With AdaBoost (T:09) 0.437 0.561 0.439 1.577 0.041
With AdaBoost (T:10) 0.432 0.561 0.439 1.289 0.071
Model Name GPT-2 Medium Qwen1.5-1.8B
Baseline 0.388 0.760 0.435 1.511 0.046
With AdaBoost (T:02) 0.448 0.760 0.477 1.571 0.041
With AdaBoost (T:03) 0.426 0.760 0.462 1.483 0.049
With AdaBoost (T:04) 0.454 0.760 0.473 1.601 0.039
With AdaBoost (T:05) 0.448 0.760 0.471 1.334 0.065
With AdaBoost (T:06) 0.465 0.760 0.470 1.249 0.076
With AdaBoost (T:07) 0.449 0.760 0.469 1.460 0.051
With AdaBoost (T:08) 0.461 0.760 0.480 1.646 0.036
With AdaBoost (T:09) 0.449 0.760 0.476 1.453 0.052
With AdaBoost (T:10) 0.447 0.760 0.483 1.154 0.090
Model Name GPT-2 Large GPT-2 XL
Baseline 0.454 0.561 0.453 2.981 0.003
With AdaBoost (T:02) 0.451 0.561 0.455 1.791 0.027
With AdaBoost (T:03) 0.458 0.561 0.451 1.954 0.020
With AdaBoost (T:04) 0.463 0.561 0.447 2.220 0.012
With AdaBoost (T:05) 0.471 0.561 0.452 2.145 0.014
With AdaBoost (T:06) 0.465 0.561 0.458 1.745 0.030
With AdaBoost (T:07) 0.459 0.561 0.453 1.729 0.031
With AdaBoost (T:08) 0.469 0.561 0.455 1.726 0.031
With AdaBoost (T:09) 0.471 0.561 0.445 1.915 0.021
With AdaBoost (T:10) 0.466 0.561 0.447 2.179 0.013
Model Name GPT-2 Large Qwen1.5-1.8B
Baseline 0.439 0.760 0.476 2.745 0.004
With AdaBoost (T:02) 0.437 0.760 0.467 1.747 0.029
With AdaBoost (T:03) 0.443 0.760 0.469 1.874 0.023
With AdaBoost (T:04) 0.445 0.760 0.460 2.018 0.017
With AdaBoost (T:05) 0.448 0.760 0.468 2.063 0.016
With AdaBoost (T:06) 0.449 0.760 0.467 1.639 0.036
With AdaBoost (T:07) 0.444 0.760 0.457 1.673 0.034
With AdaBoost (T:08) 0.453 0.760 0.468 1.727 0.031
With AdaBoost (T:09) 0.443 0.760 0.475 2.049 0.016
With AdaBoost (T:10) 0.459 0.760 0.484 2.217 0.012
Model Name GPT-2 XL Qwen1.5-1.8B
Baseline 0.463 0.763 0.504 1.165 0.089
With AdaBoost (T:02) 0.475 0.763 0.508 1.156 0.090
With AdaBoost (T:03) 0.481 0.763 0.512 0.941 0.132
With AdaBoost (T:04) 0.488 0.763 0.500 0.841 0.157
With AdaBoost (T:05) 0.481 0.763 0.518 0.821 0.162
With AdaBoost (T:06) 0.494 0.763 0.514 0.776 0.175
With AdaBoost (T:07) 0.483 0.763 0.499 0.801 0.168
With AdaBoost (T:08) 0.489 0.763 0.513 0.687 0.202
With AdaBoost (T:09) 0.492 0.763 0.516 0.832 0.159
With AdaBoost (T:10) 0.481 0.763 0.519 0.636 0.219
Table 2: This table shows weak to strong generalization using easy and hard data-splits for sciq dataset. We also study the impact of using ensemble learning methods like AdaBoost, which combines weak learners, for weak to strong training. Each model is trained for 3 epochs and uses an optimized learning rate.
Appendix F Generative Task Details
F.1 Different rating for all the datasets

### v2

@@ E.2 Important Notations
E.3 Intuition Behind Prior Term in EmsemW2S
The calculation of $\alpha$ cannot rely solely on error, $\epsilon$, as the traditional Adaboost method is valid only when $\epsilon<0.5$. Applying the same equation in our context could yield negative $\alpha$ values. We introduce a prior term, $\log(\frac{1}{1-\epsilon_{pre}}-1)$. Below we provide justification and proof regarding exponential decay of the training error using the new prior term.
Existing works on multi-class classification Adaboost Hastie et al. (2009) suggest using $\frac{1}{c}$ (where c is the number of classes) in the prior term, $\log(c-1)$, as $\frac{1}{c}$ represents the random performance of the model. However, when c (the number of classes) becomes very large, the $\log(c-1)$ term also grows significantly, causing the $\alpha$ parameters of Adaboost to become nearly identical and, consequently, less useful. To address this, we introduce a pre-training error term, $\epsilon_{pre}$, which represents an upper bound on the sample error. We then use $1-\epsilon_{pre}$ (a lower bound on accuracy) as a replacement for the $\frac{1}{c}$ term, as our model’s lowest possible accuracy is $1-\epsilon_{pre}$, not $\frac{1}{c}$.
Appendix F [Theorem 1:] Training Error Bound Exponential Decay.
Training Error of the combined weak learner is defined as the following:
$\text{Err}(H)=\frac{1}{n}\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\sum_{t=1}^{T}\alpha_{t}\mathbbm{1}(h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j})$
where $T$ is EmsemW2S rounds, $n=\sum_{i=1}^{m}k_{i}$, $k_{i}$ is the token length in the $i^{th}$ easy example $S^{e}=\{(\bm{x}^{e}_{i},\bm{y}^{e}_{i})\}_{i=1}^{m}$.
AdaBoost minimizes an upper bound on the exponential loss rather than directly minimizing the classification error. The exponential loss for our case is:
$L=\sum_{i=1}^{m}\sum_{j=i}^{k_{i}}D_{1}(i,j)\cdot e^{\sum_{t=1}^{T}\alpha_{t}\mathbbm{1}(h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j})}$
To analyze how the training error decreases, consider the weight update rule:
$D_{t+1}(i,j)\leftarrow\frac{1}{Z_{t}}D_{t}(i,j)e^{\alpha_{t}\mathbbm{1}(h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j})}\text{ forall }i\in[m],j\in[k_{i}]$
where $Z_{t}$ is the normalization factor:
$Z_{t}=\sum_{i=1}^{m}\sum_{j=i}^{k_{i}}D_{t}(i,j)\cdot e^{\alpha_{t}\mathbbm{1}(h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j})}$
Now we want the cumulative training error after $T$ rounds to be bounded by (More details regarding the following equation is in section F.1):
$\text{Err}(H)\leq\prod_{t=1}^{T}Z_{t}^{\frac{1}{\epsilon_{pre}}}$
Now for $\text{Error}_{\text{train}}$ to exponentially decay with $t$, $Z_{t}<1$. For $Z_{t}<1$ we must have $\alpha_{t}>0$.
Thus, we want
$\log{\frac{1-\epsilon_{t}}{\epsilon_{t}}}+\log(\frac{1}{1-\epsilon_{pre}}-1)>0$
$(\frac{1-\epsilon_{t}}{\epsilon_{t}})*(\frac{1}{1-\epsilon_{pre}}-1)>1$
Solving this we get
$\epsilon_{t}<\epsilon_{pre}$
This forms the our weak learning condition. This basically means that each new weak learner (LLM) that is included in the ensemble should have error less than pretrained error. This error could also be the worst possible error below which any of the LLMs would not go.
F.1 Calculating Training Error Bound for EnsemW2S
In this section we will derive equation 5 of the previous section.
We know our models weights from previous section: $\alpha_{t}=\ln\!\bigl{(}\frac{1-\epsilon_{t}}{\epsilon_{t}}\bigr{)}+\ln(\frac{\epsilon_{pre}}{1-\epsilon_{pre}})$. We can break the equation 4 in the following form,
$Z_{t}=(1-\epsilon_{t})+\epsilon_{t}\,e^{\alpha_{t}}=(1-\epsilon_{t})(\frac{1}{1-\epsilon_{pre}})=\frac{\gamma_{t}-\epsilon_{pre}}{1-\epsilon_{pre}}.$
Potential. Define the exponential potential and multi-class margin $m_{i}$
$W_{T+1}\;=\;\frac{1}{n}\sum_{i=1}^{n}\exp\!\bigl{(}-\lambda m_{i}^{(T)}\bigr{)}$
$\quad m_{i}^{(T)}=\frac{1}{2}\!\Bigl{(}F_{\bm{y}_{i}^{e,j}}^{(T)}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})-\max_{k\neq\bm{y}_{i}^{e,j}}F_{k}^{(T)}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\Bigr{)}$
where $F_{k}^{(T)}$ is the accumulated vote after $T$ rounds and $\lambda=\frac{1}{\epsilon_{pre}}$.
Mathematically, we can therefore write $F_{k}^{(T)}$ as
$F^{(T)}_{k}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})=\sum_{t=1}^{T}\alpha_{t}\mathbf{1}\left[h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})=\bm{y}_{i}^{e,j}\right]$
In a similar way we can define the following final hypothesis,
$H(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})=\arg\max_{k\in[K]}F_{k}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})$
A short algebraic check (see Zhu et al., 2006) gives the telescoping relation: $W_{t+1}=W_{t}Z_{t}^{(K-1)/K}$. Therefore,
$W_{T+1}\;=\;\prod_{t=1}^{T}Z_{t}^{\,\lambda}.$
Bounding the indicator. If $H(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}$, then $m_{i}\leq 0$. Hence $\mathbbm{1}[H(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}]\leq e^{-m_{i}^{(T)}}$ and $e^{-m_{i}^{(T)}}\leq e^{-\lambda m_{i}^{(T)}}\;$($\lambda<1$),
$\text{Err}(H)\;=\;\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[H(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}]\;\leq\;\frac{1}{n}\sum_{i=1}^{n}e^{-\lambda m_{i}}.$
The $m_{i}$ appearing in Eq. (2) is the margin after all $T$ rounds. Therefore the right-hand side of equation 8 is exactly the uniform-average potential in (7):
$\boxed{\text{Err}(H)\;\leq\;W_{T+1}=\prod_{t=1}^{T}Z_{t}^{\frac{1}{\epsilon_{pre}}}}$
Final bound. Using $1+z\leq e^{z}\;(z\in[0,1])$ we have $Z_{t}\leq e^{K\gamma_{t}}$. Here $K=\frac{1}{1-\epsilon_{pre}}$ and $\gamma_{t}=1-\epsilon_{t}-\frac{1}{K}$
$\boxed{\;\text{Err}(H)\;\leq\;\exp\!\Bigl{(}-(K\!-\!1)\sum_{t=1}^{T}\gamma_{t}\Bigr{)}\;}$
Thus, if every $\gamma_{t}>0$ the empirical error decays exponentially with $T$.
$\square$
Appendix G Margin theory for generalization
For any $\theta>0$ define the small–margin set $\mathcal{M}(\theta)=\{i:m_{i}^{(T)}\leq\theta\}$. From the derivation above
$\Pr_{i\in S}\bigl{[}m_{i}^{(T)}\leq\theta\bigr{]}\;\leq\;e^{\lambda\theta}\,W_{T+1}\;\leq\;\exp\!\Bigl{(}\lambda\theta-(K\!-\!1)\sum_{t=1}^{T}\gamma_{t}\Bigr{)}.$
VC-style margin bound (multi-class).
Let $d\!=\!\mathrm{VCdim}(\mathcal{H})$ for the weak-learner family. A direct extension of the binary margin bound of Schapire, Freund, Bartlett & Lee (1998) to $K$ classes (see Schapire & Singer, 1999; Zhu et al., 2006) states:
Theorem G.1 (Generalisation via Margins)
For any $\theta\in(0,1]$ and $\delta\in(0,1)$, with probability $\geq 1-\delta$ over the draw of an $n$-sample $S$, the boosted classifier $H$ satisfies
$\Pr_{(x,y)\sim\mathcal{D}}\!\bigl{[}H(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\bigr{]}\;\leq\;\Pr_{i\in S}\!\bigl{[}m_{i}^{(T)}\leq\theta\bigr{]}\;+\;\sqrt{\frac{c\bigl{(}d\log n+\log(1/\delta)\bigr{)}}{n\,\lambda^{2}\theta^{2}}}\!,$
where $c$ is an absolute constant.
Combining (B.1) with Thm. G.1.
Using the empirical bound (B.1) inside the theorem yields
$\boxed{\;\Pr_{\mathcal{D}}\!\bigl{[}H(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\bigr{]}\;\leq\;\exp\!\Bigl{(}\lambda\theta-(K\!-\!1)\sum_{t=1}^{T}\gamma_{t}\Bigr{)}\;+\;\sqrt{\frac{c\bigl{(}d\log n+\log(1/\delta)\bigr{)}}{n\,\lambda^{2}\theta^{2}}}\;.}$
Choose a fixed margin target $\theta>0$. Because the first term decays exponentially in $T$ (via $\sum_{t}\gamma_{t}$) and the second shrinks as $O(n^{-1/2})$, the overall test error can be made arbitrarily small by taking enough boosting rounds and/or sufficiently many training examples.
G.1 Discussion
  • Empirical error. Theorem F.1 shows that EnsemW2S drives the training error to zero at an exponential rate provided every weak learner beats random guessing $(\epsilon_{t}<\epsilon_{pre})$.
  • Generalization. Theorem G connects this fast margin growth to a quantitative VC-style bound on test error, mirroring the classical binary AdaBoost margin theory.
  • Practical implication. In practice, once the empirical small-margin fraction $\Pr_{i\in S}[m_{i}\leq\theta]$ plateaus, additional boosting offers diminishing returns unless more data (larger $n$) or a richer weak-learner class (larger $d$) is available.
Appendix H Experimental Setup
H.1 Cross Validation method for difficulty rating generation
Easy $(\bm{x}^{e},\bm{y}^{e})$ and Hard $(\bm{x}^{h},\bm{y}^{h})$ Data Split. We generate difficulty ratings using cross-validation in $n$ folds, where the model trains on $(n-1)$ splits and tests on the remaining split, repeating for all folds. The errors aggregated across the samples serve as difficulty ratings. Low-rated samples are used for weak model training, while high-rated samples are used to generate strong model training data and testing data at random. This cross-validation approach, applied to both classification and generation tasks, ensures robust difficulty splits. Additional details and plots of difficulty ratings are provided in Figures 6, 7, and 8 in the Appendix. Additionally, the continuous difficulty labels provided by Ding et al. (2024) for math, programming, chess, and reasoning datasets enable flexible easy-hard data splits at any chosen difficulty threshold. As this aspect is independent of our core message, we leave its exploration to future work.
H.2 Difficulty rating for all the datasets